Two airplanes leave an airport at the same time. The velocity of the first plane is 720 m/h at a heading 48.5°. The velocity of the second is 570 m/h at a heading of 148°. How far apart are they after 1.9 h ?
Answer in units of m.
1
Expert's answer
2020-10-05T10:56:36-0400

Distance of the first plane "720\cdot 1.9=1368(m)"


Distance of the second plane "570\cdot 1.9=1083(m)"


The angle between them "148\u00b0-48.5\u00b0=99.5\u00b0"


Using the law of cosine "d^2=1368^2+1083^2-2\cdot1368\cdot1083\cdot \cos99.5° \to"


"d\approx1880(m)"